OpenStudy (anonymous):

If cos theta =13/12 and tan theta <0 what is csc theta

OpenStudy (raden):

cos theta =13/12 or 12/13 ??

OpenStudy (anonymous):

Cos theta 13/12

OpenStudy (anonymous):

Cos theta 12/13

OpenStudy (anonymous):

Cos theta 13/12 seems wrong .. but that is what the Q asks

OpenStudy (anonymous):

yes .. @rdihawk is correct ..

OpenStudy (raden):

we knowed the maximum value for cosine is 1, while ur equation given cos theta = 13/12 = 1.083 (THIS a BIG FALSE statement)

OpenStudy (anonymous):

yeah.. you're correct. haha! :D

OpenStudy (anonymous):

So the answer is non existent?

OpenStudy (anonymous):

are you sure it is cosine and not cosecant or cotangent? cosine cannot be larger than one

OpenStudy (anonymous):

maybe it is a typo

OpenStudy (anonymous):

Yes that is how the Q is typed. And that no side of a triangle can be longer than the HYP. I will just give the answer a great big empty set.

OpenStudy (anonymous):

Thanks!